OpenStudy (anonymous):

(4√6)/(√30) rationalizing the denominator plz show work

OpenStudy (anonymous):

\[\sqrt{30}\] multiply and divide it

OpenStudy (anonymous):

Multiply and divide by \(\sqry{30}\).

OpenStudy (anonymous):

Multiply and divide by \(\sqrt30}\)

OpenStudy (anonymous):

Never mind..

OpenStudy (anonymous):

\[\frac{4 \sqrt{6} \times \sqrt{30}}{\sqrt{30}} \implies \frac{4 \sqrt{4 \times 9 \times 5}}{30}\]

OpenStudy (anonymous):

\[\frac{ 4\sqrt{6}}{ \sqrt{30} } \times \frac{\sqrt{30} }{ \sqrt{30} }\]

OpenStudy (anonymous):

\[\frac{24 \sqrt{5} }{30 }\]

OpenStudy (anonymous):

\[\frac{ 4 }{\sqrt{5} }\]

OpenStudy (anonymous):

how did you get 24?

OpenStudy (anonymous):

\[\sqrt{9} = 3\]

OpenStudy (anonymous):

\[\sqrt{4}=2\]

OpenStudy (anonymous):

(2√5 + 3√7) can you help me with this one @rajapurna

OpenStudy (anonymous):

same number with - symbol

OpenStudy (anonymous):

(2√5 + 3√7)^2

OpenStudy (anonymous):

\[2\sqrt{5}-3\sqrt{7}\]

OpenStudy (anonymous):

to do the Square ?

OpenStudy (anonymous):

so subtract then square

OpenStudy (anonymous):

No you need to multiply with 2√5−3√7

OpenStudy (anonymous):

(a + b)(a - b) = \[a ^{2} - b^{2}

OpenStudy (anonymous):

ok thank you